OpenStudy (anonymous):

1. LeAnne leaves town traveling at an average speed of 51 mph. After 4 hours, Bart leaves town traveling in the same direction at an average speed of 75 mph. Which of the following equations could be used to represent the distance between LeAnne and Bart after x hours? (Let x represent the time in hours that Bart has been traveling and y represent the distance between LeAnne and Bart.) y = 204 - 24x y = 24x y = 204 - 75x y = 204 + 24x

OpenStudy (anonymous):

at x = 0 the distance between then is 51*4 = 204 now every hour the distance decreases by 75-51 = 24 miles so y= 204-24x